# Representation Learning of Multivariate Time Series using Attention and Adversarial Training

### v1

@@ 1.3 Problem formulation
Inspired by the work of [45] and [27] an autoencoder is a suitable candidate architecture for stabilizing GAN training. Not only does it contribute in learning a latent representation of a complex distribution, but also reduces mode collapse [37] and makes it easier to perform complex modifications, e.g. through interpolation in the latent space. Since a Transformer comprises an encoder-decoder network, it can easily be adapted into an autoencoder framework that projects the input into a lower dimensionality using a bottleneck. There have been already previous work that utilize GAN training for Transformer networks, for example in developing frameworks for univariate [40] or multivariate [45] time series forecasting. In [22] the authors used a GAN scheme solely based on Transformer encoders for time series representation learning. The goal of this work is to develop an autoencoder for multivariate time series representation learning that is based on both the encoder and decoder of a Transformer network, and organize the latent space by incorporating a GAN training scheme. This way, new plausible time series samples should be generated artificially. The procedure is accompanied by an example dataset. Afterwards the latent space is evaluated using Dynamic Time Warping (DTW), Entropy scores and t-SNE representations [26].
@@ 2.1 Model architectures
The model is based on a Transformer network, i.e. an encoder $\zeta_{T}$ and a decoder $\eta_{T}$. While the original paper serves as a reference for a detailed description of each individual component [38], here the changes that turn the model into an adversarial autoencoder for multivariate time series data are presented. The input $X\in\mathbb{R}^{l\times v}$ is a multivariate time series of length $l$ and $v$ numbers of variables and comprises a sequence of $l$ feature vectors $x_{t}\in\mathbb{R}^{v}$ for $x=[x_{1},x_{2},...,x_{l}]$. As the model operates on continuous data instead on sequences of discrete word indices, the Embedding Layer projects the feature vectors $x_{t}$ into a $d$-dimensional vector space: $u_{t}=W_{e}x_{t}+b_{e}$, where $d$ is the model dimension, $W_{e}\in\mathbb{R}^{d\times l}$, $b_{e}\in\mathbb{R}^{d}$ are learnable parameters and $u_{t}\in\mathbb{R}^{d}:U\in\mathbb{R}^{l\times d}=[u_{1},u_{2},...,u_{l}]$ represents the input that corresponds to the word vectors of the original paper. This method is proposed by [43]. Subsequently, since the Transformer architecture is insensitive to any sequential ordering of the input, a Positional Encoding Layer adds a notion of time dependence. This is done by sinusoidal encodings $W_{pos}$ as proposed in the original paper: $U^{\prime}=U+W_{pos}$, where $W_{pos}\in\mathbb{R}^{l\times d}$ contains sines and cosines of different frequencies per model dimension $d$. The Transformer encoder $\zeta_{T}$ then computes the latent memory $Z\in\mathbb{R}^{l\times d}$ using the multi-head self-attention mechanism: $Z=\zeta_{T}(U^{\prime})$. To enforce a lower representation of the latent memory, two additional Feed Forward Layers incorporate a bottleneck mechanism, in which the dimensions of $Z$ are concatenated and then compressed: $Z^{\prime}\in\mathbb{R}^{k}=W_{enc}Z+b_{enc}$, where $W_{enc}\in\mathbb{R}^{k\times(ld)}$ and $b_{enc}\in\mathbb{R}^{k}$ are learnable parameters. Considering $k\ll ld$, this principle turns the Transformer into an autoencoder. A concluding Tanh Layer scales the compressed latent memory into the interval $(-1,1)$, which gives more control over the sampling limits when shaping the distribution of the latent space [31]. For the reconstruction of the time series $X$, the second Feed Forward Layer scales the compressed latent memory $Z^{\prime}$ back to the initial dimensions $Z^{\prime\prime}\in\mathbb{R}^{l\times d}$ with learnable parameters $W_{dec}$ and $b_{dec}$ similar to the preceding layer. Since this model aims to optimize a reconstruction problem rather than a classification or regression problem, the input of the Transformer decoder consists of the same input $X$, shifted to the right by $1$ time step and denoted as $\bar{X}$, together with the latent memory $Z^{\prime\prime}$: $\hat{Y}=\eta_{T}(\bar{X},Z^{\prime\prime})$, where $\hat{Y}\in\mathbb{R}^{l\times v}$ is the reconstructed multivariate time series.
@@
The adversarial training process is incorporated into the model: The generator $G$ randomly samples a memory vector $Z$ from a uniform distribution with the interval $[-1,1)^{k}$, i.e. $Z\stackrel{{\scriptstyle iid}}{{\sim}}U[-1,1)^{k},Z\in\mathbb{R}^{k}$ and decodes the sampled memory into an artificial multivariate time series $Y\in\mathbb{R}^{l\times v}$ using the Transformer decoder $\eta_{T}$. As the decoder requires a reference time series $\bar{X}$ which does not exist for a sampled memory, the decoding procedure is done in an iterative fashion: Starting with a predefined <SOS>-Token, the time series is built up step-wise by appending the current output of the decoder to the time series until the maximum sequence length is reached or a predefined <EOS>-Token is obtained. The completed time series then corresponds to the reconstruction of the memory, i.e. a fake time series created by $G$. It is important to note that this auto-regressive, generative strategy may cause error accumulation during inference [40].
@@
To compare the performance of this model with an existing approach, a convolutional autoencoder similar to [13] was implemented as well. Here, the encoder $\zeta_{C}$ consists of $8$ 1-D Convolutional Layers and has $u\in\{2,4,8,16,\allowbreak 32,64,128,256\}$ kernels of size $s\in\{21,18,15,13,11,8,\allowbreak 5,3\}$ per layer. A subsequent Aggregation Layer applies a single kernel to all $256$ channels, whose output length is then compressed to the latent dimension $k$ using a Feed Forward Layer to achieve the same autoencoder bottleneck dimensionality. The decoder $\eta_{C}$ has a symmetric structure with the parameters in the reverse order. Since the network operates with 1-D convolutions, all variables of the multivariate input $X\in\mathbb{R}^{l\times v}$ are concatenated into a single dimension $X_{concat}\in\mathbb{R}^{1\times(lv)}$. Both models are implemented in Python using PyTorch (see Appendix 0.A).
@@ 2.2 Dataset and preprocessing
In this project a subset111timeseriesclassification.com/description.php?Dataset=NATOPS of the NATOPS dataset [36] is used as an exemplary small dataset, which contains body sensor recordings of gestures used as aircraft handling signals. The data is collected by sensors on the hands, elbows, wrists and thumbs which recorded the x,y,z coordinates relatively to the person. The gestures were originally recorded at $20$ FPS (with an average duration of $2.34$ seconds). In the given subset, all sequences are normalized to a sequence length of $51$ time steps. In total, the training and validation set each contain $180$ sequences with $24$ features (sensor recordings) and $6$ classes representing different gesture commands, evenly distributed over both datasets. Figure 1 shows a t-SNE representation of the whole validation set. The plot illustrates that three classes are very well separable as different modes, while the other three classes have a very similar inherent structure. Figure 3 in the appendix in Section 0.A.2 shows exemplary plots of the dataset.
@@ 2.3 Training schedule
The algorithm for the training schedule is given below. Instead of using the objective function of (Eq. 1), the generator is trained to maximize $\log(D(G(Z)))$, which provides stronger gradients and thereby counteracts early vanishing [12]. For simplicity, the Transformer autoencoder, which consists of the encoder $\zeta_{T}$, the decoder $\eta_{T}$, and the architectural adaptions specified in Section 2.1, is denoted as $\Psi$.
@@
    Update the autoencoder by its stochastic
  • gradient: {fleqn}
@@
    Sample minibatch of $b$ fake samples
  •              $\{Y^{(1)},...,Y^{(b)}\}$ using the generator $Y=G(Z)$
    and $Z\sim p_{noise}(Z)$
    Update the discriminator by its stochastic
  • gradient: {fleqn}
@@
    Update the generator by its stochastic gradient:
  •              {fleqn}
@@
    Given the NATOPS dataset, the Transformer autoencoder was trained for 2 00020002\,0002 000 epochs, with a batch size equals to $32$ and a learning rate of $10^{-4}$. The training was done at the Training Center for Machine Learning (TCML) Cluster in Tübingen (Grant number 01IS17054). The convolutional autoencoder was trained using the same procedure. However, all weights were instead initialized from a zero-centered Normal distribution with a standard deviation of $0.02$ [31] and a learning rate of $0.001$ was used.
    A variant of the above algorithm is yielded by incorporating the Wasserstein GAN scheme [2] to stabilize GAN training (an overview of different stabilization strategies is given in the appendix in Section 0.A.1). Here, both the discriminator and the generator are optimized using an adapted loss function (eq. (5)), the discriminator uses a linear activation instead of sigmoid, is trained $5$ times more than the generator in each iteration and its weights are constrained to a range of $[-0.1,0.1]$ after each update. Furthermore RMSprop is used as learning rule with a small learning rate of $5\cdot 10^{-5}$ and without momentum.
@@ 2.4 Evaluation metrics
    The characteristic of the latent space and the generation of artificial multivariate time series can be evaluated qualitatively and quantitatively [8]. However, unlike image-based GANs, where the Inception Score [33] or the Fréchet Inception Distance (FID) [15] are established metrics, there are no standards set for time series data, especially not for multivariate time series. [7] gives an overview of different evaluation metrics, which however are mainly focused on image generation and for the most part lack possibilities to adapt to (multivariate) time series data.
@@ 3 Results
    Both the Transformer autoencoder and the convolutional autoencoder were trained with and without the GAN training scheme. Training without GAN only involved minimizing the reconstruction error without any regularization of the latent space. Additionally, both models were trained using the Wasserstein GAN approach. In this section the models are abbreviated the following: TAE = Transformer autoencoder, TAE-GAN = Transformer autoencoder with GAN scheme,
    Figure 2: t-SNE representation of all models. The blue dots refer to the true time series from the validation dataset, while the orange dots refer to artificially generated time series from the model.
    TAE-WGAN = Transformer autoencoder with Wasserstein GAN scheme. The convolutional autoencoder is abbreviated in the same manner using the basis CAE. Figure 2 shows the t-SNE visualizations of all models. Table 1 contains the results of all models regarding similarity (Avg. DTW), diversity (Entropy) and the reconstruction ability of the original signals (Test Error).
@@
    TAE
    $28.273\pm 3.706$
    $0.544\pm 0.018$
    0.018
    TAE-GAN
    $36.918\pm 1.086$
    $0.427\pm 0.003$
    $0.283$
    TAE-WGAN
    $\textbf{19.919}\pm 3.893$
    $0.394\pm 0.032$
    $0.019$
    CAE
    $36.524\pm 2.345$
    $0.284\pm 0.005$
    $0.082$
    CAE-GAN
    $35.164\pm 2.273$
    $0.589\pm 0.007$
    $0.047$
    CAE-WGAN
    $37.949\pm 2.636$
    $\textbf{0.855}\pm 0.006$
    $0.024$
    3.0.1 TAE:
    The model shows a higher diversity (Entropy) compared to the TAE trained with GAN. The t-SNE representation in Figure 2 illustrates this variance, but the generated time series are hardly represented within the modes of the dataset, as the latent space was not shaped any further. The reconstruction error is smaller compared to the TAE trained with GAN. Apparently, the GAN training counteracted decoding accuracy as a regulatory constraint. Although TAE achieved the lowest reconstruction error, it is important to note that the Transformer model generally suffers from error accumulation when generating artificial time series signals, because the generation process is done in an interative fashion. Figure 6 in the appendix in Section 0.A.2 shows exemplary generated time series that have similar patterns to those from the dataset, but also sudden disruptions that are not represented in the dataset (Example 1 and 8). In addition, small fluctuations are contained: Straight lines and smooth curves could not be reproduced adequately.
@@ 3.0.2 TAE-GAN:
    Both the diversity (Entropy) and the similarity (DTW) are worse compared to TAE. This is also suggested by the t-SNE representation, since only one mode is learned that does not overlap with the original distribution. This is a typical example of 1. Mode dropping and 2. Mode collapse, where 1. mass is put outside of the support of the underlying distribution of the dataset and 2. only a single type of output is learned and other modes are disregarded [25]. The generated time series in Figure 9 in the appendix in Section 0.A.2 confirm the low diversity, and hardly show any similarity to the dataset. However, the error plots in Figure 8 in the appendix in Section 0.A.2 show stable training and illustrate that although the discriminator became increasingly uncertain after 2 00020002\,0002 000 epochs, overall convergence was not achieved. The generative quality could maybe have been higher if the training had lasted longer.
@@ 3.0.3 TAE-WGAN:
    With a score of $19.919$, the similarity (DTW) to the dataset is the highest compared to all other models. However, the diversity (Entropy) has deteriorated compared to TAE and TAE-GAN. The t-SNE representation in Figure 2 actually suggests a higher score due to the higher dispersion, but it also shows that the three well-separated modes are not sufficiently represented. The reconstruction ability was hardly influenced by the GAN regularization compared to TAE (similar Test Error) and is smaller compared to the convolutional models. The generated time series in Figure 11 in the appendix Section 0.A.2 have similar patterns to those from the dataset and indeed show some diversity. However, they also contain fluctuations where straight lines or smooth curves were expected. The error plots in Figure 10 in the appendix in Section 0.A.2 illustrate stable training and indicate that convergence was achieved.

### v2

@@ 1.3 Problem Formulation
Inspired by the work of [38] and [24] an autoencoder is a suitable candidate architecture for stabilizing GAN training. Not only does it contribute in learning a latent representation of a complex distribution, but also reduces mode collapse [32] and makes it easier to perform complex modifications, e.g. through interpolation in the latent space. Since a Transformer comprises an encoder-decoder network, it can easily be adapted into an autoencoder framework that projects the input into a lower dimensionality using a bottleneck. There have been already previous work that utilize GAN training for Transformer networks, for example in developing frameworks for univariate [34] or multivariate [38] time series forecasting. In [20] the authors used a GAN scheme solely based on Transformer encoders for time series representation learning. The goal of this work is to develop an autoencoder for multivariate time series representation learning that is based on both the encoder and decoder of a Transformer network, and organize the latent space by incorporating a GAN training scheme. This way, new plausible time series samples should be generated artificially. The procedure is accompanied by an example dataset. Afterwards the latent space is evaluated using Dynamic Time Warping (DTW), Entropy scores, and t-SNE representations [23].
@@ 2.1 Model Architectures
The model is based on a Transformer network, i.e. an encoder $\zeta_{T}$ and a decoder $\eta_{T}$. While the original paper serves as a reference for a detailed description of each individual component [33], here the changes that turn the model into an adversarial autoencoder for multivariate time series data are presented. The input $X\in\mathbb{R}^{l\times v}$ is a multivariate time series of length $l$ and $v$ numbers of variables and thus comprises a sequence of $l$ feature vectors $x_{t}\in\mathbb{R}^{v}$ for $x=[x_{1},x_{2},...,x_{l}]$. As the model operates on continuous data instead on sequences of discrete word indices, the Embedding Layer projects the feature vectors $x_{t}$ into a $d$-dimensional vector space: $u_{t}=W_{e}x_{t}+b_{e}$, where $d$ is the model dimension, $W_{e}\in\mathbb{R}^{d\times v}$, $b_{e}\in\mathbb{R}^{d}$ are learnable parameters and $u_{t}\in\mathbb{R}^{d}:U\in\mathbb{R}^{l\times d}=[u_{1},u_{2},...,u_{l}]$ represents the input that corresponds to the word vectors of the original paper. This method is proposed by [36]. Subsequently, since the Transformer architecture is insensitive to any sequential ordering of the input, a Positional Encoding Layer adds a notion of time dependence. This is done by sinusoidal encodings $W_{pos}$ as proposed in the original paper: $U^{\prime}=U+W_{pos}$, where $W_{pos}\in\mathbb{R}^{l\times d}$ contains sines and cosines of different frequencies per model dimension $d$. The Transformer encoder $\zeta_{T}$ then computes the latent memory $Z\in\mathbb{R}^{l\times d}$ using the multi-head self-attention mechanism: $Z=\zeta_{T}(U^{\prime})$. To enforce a lower representation of the latent memory, two additional Feed Forward Layers incorporate a bottleneck mechanism, in which the dimensions of $Z$ are concatenated and then compressed: $Z^{\prime}\in\mathbb{R}^{k}=W_{enc}Z+b_{enc}$, where $W_{enc}\in\mathbb{R}^{k\times(ld)}$ and $b_{enc}\in\mathbb{R}^{k}$ are learnable parameters. Considering $k\ll ld$, this principle turns the Transformer into an autoencoder. A concluding Tanh Layer after the first Feed Forward Layer scales the compressed latent memory into the interval $(-1,1)$, which gives more control over the sampling limits when shaping the distribution of the latent space [27]. For the reconstruction of the time series $X$, the second Feed Forward Layer scales the compressed latent memory $Z^{\prime}$ back to the initial dimensions $Z^{\prime\prime}\in\mathbb{R}^{l\times d}$ with learnable parameters $W_{dec}$ and $b_{dec}$ similar to the preceding layer, which serves as the first input for the Transformer decoder $\eta_{T}$. Since this model aims to optimize a reconstruction problem rather than a classification or regression problem, the second input of the Transformer decoder consists of the same input $X$, shifted to the right by $1$ time step and denoted as $\bar{X}$. The output of the Transformer decoder corresponds to $\hat{Y}=\eta_{T}(\bar{X},Z^{\prime\prime})$, where $\hat{Y}\in\mathbb{R}^{l\times v}$ is the reconstructed multivariate time series.
Figure 1: The model augments the Transformer architecture with an autoencoder and an adversarial training scheme. For the Transformer decoder, the input time series is shifted to the right to serve as the output signal that is to be predicted. A memory vector is drawn from a prior distribution and is decoded to generate an artificial time series. The discriminator decides if a time series is true (from the dataset) or fake (artificially generated).
@@
The adversarial training process is incorporated into the model: The generator $G$ randomly samples a memory vector $Z$ from a uniform prior distribution with the interval $[-1,1)^{k}$, i.e. $Z\stackrel{{\scriptstyle iid}}{{\sim}}U[-1,1)^{k},Z\in\mathbb{R}^{k}$ and decodes the sampled memory into an artificial multivariate time series $Y\in\mathbb{R}^{l\times v}$ using the decoder of the autoencoder and the Transformer decoder $\eta_{T}$. As $\eta_{T}$ requires a reference time series $\bar{X}$ which does not exist for a sampled memory, the decoding procedure is done in an iterative fashion: Starting with a predefined <SOS>-Token, the time series is built up step-wise by appending the current output of the decoder to the time series until the maximum sequence length is reached or a predefined <EOS>-Token is obtained. The completed time series then corresponds to the reconstruction of the memory, i.e. a fake time series created by $G$. It is important to note that this auto-regressive, generative strategy may cause error accumulation during inference [34].
@@
To compare the performance of this model with an existing approach, a convolutional autoencoder similar to [12] was implemented as well. Here, the encoder $\zeta_{C}$ consists of $8$ 1-D Convolutional Layers and has $u\in\{2,4,8,16,\allowbreak 32,64,128,256\}$ kernels of size $s\in\{21,18,15,13,11,8,\allowbreak 5,3\}$ per layer. A subsequent Aggregation Layer applies a single kernel to all $256$ channels, whose output length is then compressed to the latent dimension $k$ using a Feed Forward Layer to achieve the same autoencoder bottleneck dimensionality. The decoder $\eta_{C}$ has a symmetric structure with the parameters in the reverse order. Since the network operates with 1-D convolutions, all variables of the multivariate input $X\in\mathbb{R}^{l\times v}$ are concatenated into a single dimension $X_{concat}\in\mathbb{R}^{1\times(lv)}$. Both models are implemented in Python using PyTorch.
@@ 2.2 Dataset and preprocessing
In this work a subset111timeseriesclassification.com/description.php?Dataset=NATOPS of the NATOPS dataset [31] is used as an exemplary small dataset, which contains body sensor recordings of gestures used as aircraft handling signals. The data is collected by sensors on the hands, elbows, wrists and thumbs which recorded the x,y,z coordinates relatively to the person. The gestures were originally recorded at $20$ FPS (with an average duration of $2.34$ seconds). In the given subset, all sequences are normalized to a sequence length of $51$ time steps. In total, the training and validation set each contain $180$ sequences with $24$ features (sensor recordings) and $6$ classes representing different gesture commands, evenly distributed over both datasets. Figure 4 shows ten exemplary time series. Figure 2 shows a t-SNE representation of the whole validation set. The plot illustrates that three classes are very well separable as different modes, while the other three classes have a very similar inherent structure.
@@ 2.3 Training schedule
The algorithm for the training schedule is given above. Instead of using the objective function of (Eq. 1), the generator is trained to maximize $\log(D(G(Z)))$, which provides stronger gradients and thereby counteracts early vanishing [11]. For simplicity, the Transformer autoencoder, which consists of the encoder $\zeta_{T}$, the decoder $\eta_{T}$, and the architectural adaptions specified in Section 2.1, is denoted as $\Psi$.
@@
  • Update the autoencoder by its stochastic gradient: {fleqn}
@@
    Sample minibatch of $b$ fake samples $\{Y^{(1)},...,Y^{(b)}\}$
  • using the generator $Y=G(Z)$ and $Z\sim p_{noise}(Z)$
  • Update the discriminator by its stochastic gradient: {fleqn}
@@
    Update the generator by its stochastic gradient: {fleqn}
@@
    Given the NATOPS dataset, the Transformer autoencoder was trained for 2 00020002\,0002 000 epochs, with a batch size equals to $32$ and a learning rate of $10^{-4}$. The training was pursued at the Training Center for Machine Learning (TCML) Cluster in Tübingen (Grant number 01IS17054). The convolutional autoencoder was trained using the same procedure. However, all weights were instead initialized from a zero-centered Normal distribution with a standard deviation of $0.02$ [27] and a learning rate of $0.001$ was used.
    A variant of the above algorithm is yielded by incorporating the Wasserstein GAN scheme [1] to stabilize GAN training. Here, both the discriminator and the generator are optimized using an adapted loss function (eq. (5)), the discriminator uses a linear activation instead of sigmoid, is trained $5$ times more than the generator in each iteration and its weights are constrained to a range of $[-0.1,0.1]$ after each update. Furthermore RMSprop is used as learning rule with a small learning rate of $5\cdot 10^{-5}$ and without momentum.
@@ 2.4 Evaluation metrics
    The characteristic of the latent space and the generation of artificial multivariate time series can be evaluated qualitatively and quantitatively [7]. However, unlike image-based GANs, where the Inception Score [28] or the Fréchet Inception Distance (FID) [14] are established metrics, there are no standards set for time series data, especially not for multivariate time series. Borji in [6] gives an overview of different evaluation metrics, which however are mainly focused on image generation and for the most part lack possibilities to adapt to (multivariate) time series data.
@@ 3 Results
    Both the Transformer autoencoder and the convolutional autoencoder were train-ed with and without the GAN training scheme. Training without GAN only involved minimizing the reconstruction error without any regularization of the latent space. Additionally, both models were trained using the Wasserstein GAN approach. In this section the models are abbreviated the following: TAE = Transformer autoencoder, TAE-GAN = Transformer autoencoder with GAN scheme, TAE-WGAN = Transformer autoencoder with Wasserstein GAN scheme. The convolutional autoencoder is abbreviated in the same manner using the basis CAE. Figure 3 shows the t-SNE visualizations of all models. Table 1 contains the results of all models regarding similarity (Avg. DTW), diversity (Entropy) and the reconstruction ability of the original signals (Test Error).
    3.0.1 TAE:
    The model shows a higher diversity (Entropy) compared to the TAE trained with GAN. The t-SNE representation in Figure 3 illustrates this variance, but the generated time series are hardly represented within the modes of the dataset, as the latent space was not shaped any further. The reconstruction error is smaller compared to the TAE trained with GAN. Apparently, the GAN training counteracted decoding accuracy as a regulatory constraint. Although TAE achieved the lowest reconstruction error, it is important to note that the Transformer model generally suffers from error accumulation when generating artificial time series signals, because the generation process is done in an iterative fashion.
    Figure 3: t-SNE representation of all models. The blue dots refer to the true time series from the validation dataset, while the orange dots refer to artificially generated time series from the models. Models without GAN regularization (TAE, CAE) did not learn the representation. TAE-GAN seems to have suffered from mode dropping and mode collapse.
@@ 3.0.1 TAE:
    TAE $28.273$ $0.544$ 0.018
    TAE-GAN $36.918$ $0.427$ $0.283$
    TAE-WGAN 19.919 $0.394$ $0.019$
    CAE $36.524$ $0.284$ $0.082$
    CAE-GAN $35.164$ $0.589$ $0.047$
    CAE-WGAN $37.949$ 0.855 $0.024$
@@ 3.0.2 TAE-GAN:
    Both the diversity (Entropy) and the similarity (DTW) are worse compared to TAE. This is also suggested by the t-SNE representation, since only one mode is learned that does not overlap with the original distribution. This is a typical example of 1. Mode dropping and 2. Mode collapse, where 1. mass is put outside of the support of the underlying distribution of the dataset and 2. only a single type of output is learned and other modes are disregarded [22].
@@ 3.0.3 TAE-WGAN:
    With a score of $19.919$, the similarity (DTW) to the dataset is the highest compared to all other models. However, the diversity (Entropy) has deteriorated compared to TAE and TAE-GAN. The t-SNE representation in Figure 3 actually suggests a higher score due to the higher dispersion, but it also shows that the three well-separated modes are not sufficiently represented. The reconstruction ability was hardly influenced by the GAN regularization compared to TAE (similar Test Error) and is smaller compared to the convolutional models. The generated time series in Figure 5 have similar patterns to those from the dataset and indeed show some diversity. However, they contain fluctuations where straight lines or smooth curves were expected.
    Figure 4: Ten random examples of time series signals from the NATOPS validation set.
    Figure 5: Ten different time series samples generated by TAE-WGAN. They show similar patterns to the validation dataset, however contain small fluctuations.